1. If the mean is 40 an the standard deviation is 8, what percent of data is between 32-48?
2. If the mean is 40 and the standard deviation is 6, what number is two standard deviation below the mean?

Answer by robertb(5830)

By the empirical rule, 68% of the data is between 32-48 (1 standard deviation from the mean.)

40 - 2*6 = 40 - 12 = 28.